OpenStudy (anonymous):

solve differential equation below.... y'+2y+1/s*L(f)=sin(t) where y(0)=0 i manage to get till y=(1/s^2+1)(s/s^2+1) can someone confirm is correct ?? thanks a lot :)

OpenStudy (anonymous):

can I get some brackets

OpenStudy (turingtest):

what is L(f) ?

OpenStudy (turingtest):

laplace?

OpenStudy (turingtest):

ok yes, it's laplace, but your problem confuses me...

OpenStudy (anonymous):

yup turingtest.....that clearer...

OpenStudy (anonymous):

cuddlepony i attached a file on my post that will be easier to read the question...ty

OpenStudy (turingtest):

oh wow, frankly I've never seen a DE that had a Laplace already in it I'll be curious to see the solution too, sorry I doubt I can help I'll try though :D

OpenStudy (anonymous):

i see....but i check my notes that is equivalent to 1/s L(f) .....maybe we assume that correct...haha

OpenStudy (anonymous):

@ash2326 can you try to solve this question for me ty....??

OpenStudy (ash2326):

Let me try @angalc557

OpenStudy (ash2326):

@angalc557 Is this your question? \[(y'+2y+1) L(t)=\sin (t)\] I'm sure there can't be 1/s or L(s) before taking Laplace Transform!!!

OpenStudy (anonymous):

hmm the you can refer to the pic i attached to this post for clearer view....ty

OpenStudy (ash2326):

Ok we have \[ y'+y +\int _{0} ^{t} y(\tau) d\tau= \sin t \] Let's take Laplace tranform of this \[ sy(s)-y(0)+y(s) +\frac{1}{s} y(s)= \frac{1}{s^2+1} \] y(0)=0 so \[ y(s)(s+1+\frac{1}{s}) =\frac{1}{s^2+1}\] we get \[y(s)( \frac{s^2+s+1}{s})=\frac{1}{s^2+1}\] \[y(s)= \frac{s}{(s^2+s+1)(s^2+1)}\] s can be written as \[s=(s^2+s+1)-(s^2+1)\] so we have \[y(s)= \frac{(s^2+s+1)-(s^2+1)}{(s^2+s+1)(s^2+1)}\] now we get \[y(s)=\frac{1}{(s^2+1)}-\frac{1}{(s^2+s+1)}\] now \[s^2+s+1=s^2+s+1/4-1/4+1=(s+1/2)^2-(\frac{\sqrt{3}}{2})^2\] so we have now \[y(s)=\frac{1}{(s^2+1)}-\frac{1}{(s+1/2)^2-(\frac{\sqrt{3}}{2})^2}\] Now take inverse Laplace Transform we get \[\large{y(t)= \sin t- e^{-\frac{1}{2}t} \times \sin(\frac{\sqrt 3}{2} t)}\]

OpenStudy (ash2326):

Sorry it 'd be \[y(s)= \frac{1}{s^2+1}-\frac{1}{(s+1/2)^2+(\frac{\sqrt{3}}{2})^2}\] so we'll get \[\large{y(t)=\sin t -e^{-\frac{1}{2} t} sin(\frac{\sqrt3}{2} t)}\]

OpenStudy (ash2326):

Sorry I mistakenly put a minus sign instead of a plus. Did you get it though?

OpenStudy (anonymous):

i thought hmm 2y is the question not y ?

OpenStudy (ash2326):

Oh no:( Can you do it following the same method or should I do it again??

OpenStudy (anonymous):

y(s)(s+2+1s)=1s2+1 the middle one become two right? just confirm for me?

OpenStudy (ash2326):

Yeah , keep going

OpenStudy (anonymous):

is it y=s/(s^2+1)(s/s^2+1)

OpenStudy (anonymous):

first one is 1/(s^2+1)(s/s^2+1) right for y(s)=?

OpenStudy (ash2326):

We have \[y(s)[s+2+ \frac{1}{s}]= \frac{1}{s^2+1}\] we get \[ y(s)[\frac{s^2+2s+1}{s}]=\frac{1}{s^2+1}\] so we get \[ y(s)= \frac{s}{(s^2+2s+1)(s^2+1)}\]

OpenStudy (anonymous):

i see alright agreed...then the part you said s can be written as how u get that ? the next step?

OpenStudy (ash2326):

Now \[s= \frac{1}{2}((s^2+2s+1)-(s^2+1))\] We have \[ y(s)= \frac{(s^2+2s+1)-(s^2+1)}{2(s^2+2s+1)(s^2+1)}\] we get now \[ y(s)= \frac{1}{2(s^2+1)}-\frac{1}{2(s^2+2s+1)}\] we know \[(s^2+2s+1)=(s+1)^2\] so we get \[ y(s)= \frac{1}{2(s^2+1)}-\frac{1}{2(s+1)^2}\] Now taking inverse Laplace Transform we get \[ y(t)= \frac{1}{2}\sin t-\frac{1}{2}e^{-t} t\] Using the following \[ LT[t]=\frac{1}{s^2}\] and \[LT[e^{-t}t]=\frac{1}{(s+1)^2}\]

OpenStudy (ash2326):

@angalc557 Did you understand it?

OpenStudy (anonymous):

hmm maybe i not so good i cant figure out how u get the half there where it came from? could u explain sorry....

OpenStudy (ash2326):

No need to be sorry @angalc557 we have s in the numerator I want to write it as a difference of the terms in denominator If I write \((s^2+2s+1)-(s^2+1)\) I'll get 2s to get s, I have divided it by 2 \[\huge{\frac{(s^2+2s+1)-(s^2+1)}{2}=s}\]

OpenStudy (anonymous):

y(t)=1/2 sin t - 1/2 te^-t right? i understand adi till the end.....ty a lot :)